Upon examining the contents of 38 backpacks, it was found that 23 contained a black pen, 27 contained a blue pen, and 21 contained a pencil, 15 contained both a black pen and a blue pen, 12 contained both a black pen and a pencil, 18 contained both a blue pen and a pencil, and 10 contained all three items. How many backpacks contained none of the three writing instruments?
Answer by Edwin McCravy(20056)



There are 3 overlapping sets A,B, and C inside the big rectangle,
making 8 regions to consider.   

Circle A contains all the backpacks that had at least a black pen.
Circle B contains all the backpacks that had at least a blue pen.
Circle C contains all the backpacks that had at least a pencil.

We want to know how many are in
region k, which is the region outside all three circles.

Let's reverse the order the clues
are listed in:

10 contained all three items.
So we put 10 in region f



18 contained both a blue pen and a pencil,
We have accounted for 10 of these among the 18, so the remaining
8 are in region i.  So we put 8 in i 



12 contained both a black pen and a pencil,
We have accounted for 10 of these among the 12, so the remaining
2 are in region g.  So we put 2 in g



15 contained both a black pen and a blue pen,
We have accounted for 10 of these among the 15, so the remaining
5 are in region e.  So we put 5 in region e:



21 contained a pencil,
We have filled in 3 of the regions of the bottom circle 
representing all the backpacks which had at least a pencil.
2+10+8=20, so the remaining 1 of the 21 is in region j.
So we put 1 in region j.



27 contained a blue pen

We have filled in 3 of the regions of the right circle
representing all the backpacks which had at least a blue pen.
5+10+8=23, so the remaining 4 of the 27 is in region h. So
we put 4 in region h.


23 contained a black pen,
We have filled in 3 of the regions of the left circle 
representing all the backpacks which had at least a black pen.
5+10+2=17, so the remaining 6 of the 23 is in region h.  So
we put 6 in region d.



Now we come to the question:

How many backpacks contained none of the three writing instruments?
Now we have the number in every region but k.  We are told
there are 38 backpacks.  The other 7 regions contain

10+8+2+5+1+4+6 = 36

So the remining 2 are in the outside region k.  So we
put 2 in region k:



Answer: 2

Edwin